# A Calculus of Inheritance

### v4

@@ 6.1. Emergent Phenomena
Several features of inheritance-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
Self-referential records with lazy evaluation:
A mixin’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
Open recursion with symmetric inheritance:
Inheritance merges definitions from independent sources, all sharing the same inherited mixin tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
Commutative, idempotent merge:
The inheritance operation is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
@@
Every inheritance-calculus operation is defined by recursion on mixin trees, not on individual values. When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued mixins $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
@@
In a web application (Yang, 2026) initially targeting Python’s synchronous sqlite3 and http.server modules, the business logic MIXINv2 file contained no async-specific code. When the FFI layer was replaced with one built on aiosqlite and starlette, the business logic MIXINv2 file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the MIXINv2 file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
Immunity to nonextensibility:
Programming languages typically provide a mix of extensible and nonextensible constructs. In such languages, extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d. S. Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d. S. Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework. In contrast, every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3). There is no nonextensible construct, so it is impossible to write a nonextensible program. In the Case Study (Section 5), new operations and data types are added without modifying existing files. Tree-level inheritance (Section 5) and multi-path qualified this (equation 6; Appendix B) account for this. The Expressive Asymmetry (Section 4.2; Theorem A.3) provides the formal counterpart: the mixin trees in its separating construction are mechanical translations of $\lambda$-terms via $\mathcal{T}$ (Section 4), yet a new observable projection is added to them without modification. This is the extensibility that the Expression Problem (Wadler, 1998) demands. [^3] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement. A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.
Each of these features is primitive in inheritance-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms. In common sense, inheritance-calculus is strictly more expressive than the $\lambda$-calculus.
@@ 6.2. Semantic Alternatives
Every language whose programs are tree-structured ASTs with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another. The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
@@
We now classify the design space to explain why we chose the semantics of Section 3. Two axes remain free: the composition mechanism—how the single reference type incorporates definitions; and the reference resolution—how the single reference type finds its target.
@@
Conditional breaks monotonicity
@@
Same-label definitions are rejected as errors. All three identities hold—but vacuously: the identities are satisfied because the compositions that would test them are refused. Conflict rejection prevents extending existing labels (Harper and Pierce, 1991)—the composability that the Expression Problem (Wadler, 1998) requires.
@@
$A+B\neq B+A$: one source takes precedence (Cunningham, 2014; Dolstra et al., 2010), introducing an ordering dependency.
@@
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. This applies to every language whose composition mechanism is function application—functional, object-oriented, and multi-paradigm alike—because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction. The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in. The frameworks that simulate universal extensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
@@
Removing or conditionally including definitions breaks monotonicity, so the semantic equations (Appendix C) may have no fixed point (Tarski, 1955).
Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and universally extensible. The other candidates each introduce a vulnerability to nonextensibility or break monotonicity.
@@
With dynamic scope, references resolve at the point of use rather than the definition site. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Dynamic scope breaks the $\lambda$-calculus embedding (Section 4). The Substitution Lemma (Lemma D.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect. With dynamic scope this separation fails. [^4] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$. The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$). With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$. With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$. This is the classical variable-capture failure of dynamic scoping. As with early binding, the mixin cannot serve as a complete replacement for functions.
@@
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (D.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix B). This rejects nested extensions where two independently authored modules define the same option—a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
Conditional breaks monotonicity
@@
The semantics of inheritance-calculus was not chosen from a menu. We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands (three successive implementations confirmed this—see above). Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@ 7. Related Work
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
@@
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991), and a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem; Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975). Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions). In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies. None of these approaches yields a semantics of the lazy $\lambda$-calculus that simultaneously lives in naive set theory and is fully abstract: Scott’s $D_{\infty}$ is adequate but not fully abstract, and game semantics is fully abstract but requires machinery (arenas, strategies, innocence conditions) far beyond set theory. None of them eliminates the function space from the semantic domain; none avoids capture-avoiding substitution; and none dispenses with reduction rules or evaluation strategies.
@@
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. Graph models (Engeler, 1981; Plotkin, 1993) take a web $A$ equipped with an injection $\iota:\mathcal{P}_{\mathrm{fin}}(A)\times A\hookrightarrow A$ and set $D=\mathcal{P}(A)$. Application is $f\cdot x=\{\,b\mid\exists\,S\subseteq x,\;\iota(S,b)\in f\,\}$: the function space $D\to D]$ is not constructed explicitly but encoded inside $D$ via $\iota$. The construction uses only powerset lattices, no Scott topology or continuity conditions. Graph models are sensible (they equate all unsolvable terms) but not fully abstract: Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$, which is strictly coarser than observational equivalence for the lazy $\lambda$-calculus. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the filter (upward-closed, intersection-closed set) of all intersection types derivable for it. The collection of all such filters forms an $\omega$-algebraic lattice. Barendregt, Coppo, and Dezani-Ciancaglini proved that type assignment is sound and complete with respect to the filter $\lambda$-model. Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the $\lambda$-theory of the BCD filter model coincides exactly with Böhm tree equality: two terms receive the same set of types if and only if they have the same Böhm tree. Filter models are more discriminating than graph models but are still not fully abstract for the lazy $\lambda$-calculus, and require the additional machinery of intersection type theory and algebraic lattices. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic: the exponential modality !! is interpreted via finite multisets $\mathcal{M}_{\mathrm{fin}}(A)$, giving a co-Kleisli category whose morphisms are relations from $\mathcal{M}_{\mathrm{fin}}(A)$ to $B$. The multiset structure corresponds to non-idempotent intersection types (de Carvalho, 2018), which carry quantitative information (bounds on reduction length) that idempotent systems lack. Ehrhard (Ehrhard, 2005) showed that in the finiteness-space refinement of this model, standard fixed-point operators are not finitary, and a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. For the lazy $\lambda$-calculus, full abstraction has been achieved only through game semantics (Abramsky and Ong, 1993; Abramsky and McCusker, 1995); none of the three set-theoretic families attains it. All three families are semantic models of the $\lambda$-calculus: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure. None of them is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds. Moreover, all three identify terms more coarsely than observational equivalence—graph and filter models coincide with Böhm tree equality, while relational models cannot even host a standard reflexive object (Ehrhard, 2005)—and none provides a fully abstract semantics for the lazy $\lambda$-calculus without resorting to game-theoretic machinery.

### v6

@@ 6.1. Emergent Phenomena
In developing real programs in MIXINv2, we encountered three phenomena that we did not design for. They can be seen in the source code (Yang, 2026) and the Case Study (Section 5). Below we explain why they arise from the three primitives.
@@
Single-valued computation and set-valued computation are the same mechanism. The semantic equations of Section 3 resolve every property to a set of paths, so operations defined on single values automatically distribute over all combinations of set-valued inputs (Section 5). When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise. [^9] Inheritance-calculus to relational algebra: path from root to leaf == tuple; trie union == relation (set of tuples); $\mathtt{bases}$ (equation 4) == cross-join; $\mathtt{supers}$ (equation 2) == union; $\mathtt{properties}$ (equation 1) == projection. The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths. This contrasts with Datalog’s bottom-up materialization, but the relational structure of the computation is the same. Appendix F makes this correspondence precise with a systematic encoding of Datalog into inheritance-calculus, demonstrating that variable joins, constant matches, multi-body rules, and recursive rules are all expressible using only the three primitives of Section 2.
Immunity to nonextensibility:
It is structurally impossible to write a nonextensible program. Every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3). The Case Study (Section 5) and the Expressive Asymmetry (Section 4.2; Theorem C.7) demonstrate and formalize this: new operations and data types are added, and new observable projections are introduced. This is the extensibility that the Expression Problem (Wadler, 1998) demands. [^10] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement. A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs. Programming languages typically provide a mix of extensible and nonextensible constructs, where extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d. S. Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d. S. Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework. Section 6.2 analyzes why deep merge is the only composition mechanism that achieves this.
@@
Business logic in MIXINv2 declares abstract property slots; different assemblies fill them with synchronous or asynchronous implementations via inheritance. Because the slot’s color varies at assembly time without affecting the code that uses it, the same MIXINv2 file runs under both synchronous and asynchronous runtimes (Yang, 2026), an instance of function color blindness (Nystrom, 2015). The mechanism is dependency injection via inheritance: business logic declares empty slots, and the assembly inherits an FFI module that fills them, so the same business logic accepts implementations of any color. [^11] Function color (Nystrom, 2015) arises when effectful functions have signatures $a\to f\,b$ (Kleisli arrows), requiring monadic bind to compose and making the effect $f$ visible at every call site. Because inheritance-calculus composes program structure rather than executing effects, this type distinction does not arise at the calculus level. The color is determined at assembly time by the choice of inherited implementation, and effect execution is delegated to the host language. Unstructured control flow, such as capturing and resuming continuations as in call/cc, cannot be abstracted this way, as it requires a global CPS transformation that inheritance-calculus does not express.
@@ 6.2. Semantic Alternatives
The analysis below assumes that the AST is a named tree:every edge from parent to child carries a label, so that every node is uniquely identified by the path from the root. For configuration languages this is naturally the case: every key in a JSON, YAML, or Nix attribute set is a label, and nesting produces paths. [^12] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL. The NixOS module system (NixOS Contributors, 2024), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree. For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels. After ANF transformation (Flanagan et al., 1993), however, every intermediate result is bound to a name, and the remaining anonymous positions [^13] For example, the body of a $\mathbf{let}$. can be assigned synthetic fresh labels such as $\mathrm{result}$ and $\mathrm{tailCall}$ in Section 4, yielding a named tree.
Under this premise, every language whose programs are such named trees with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another. The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
@@
We now classify the design space to explain why we chose the semantics of Section 3. Two axes remain free: the composition mechanism, which determines how the single reference type incorporates definitions; and the reference resolution, which determines how the single reference type finds its target.
@@
Conditional causes divergence
@@
Same-label definitions are rejected as errors. All three identities hold, but vacuously: the identities are satisfied because the compositions that would test them are refused. Conflict rejection prevents extending existing labels (Harper and Pierce, 1991), precisely the composability that the Expression Problem (Wadler, 1998) requires.
@@
One source takes precedence (Cunningham, 2014; Dolstra et al., 2010), so $A+B\neq B+A$, introducing an ordering dependency.
@@
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. This applies to every language whose composition mechanism is function application, whether functional, object-oriented, or multi-paradigm, because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction. The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in. The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
@@
Removing or conditionally including definitions can introduce cycles in the dependency graph, causing the recursive evaluation (Appendix A) to diverge.
Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and immune to nonextensibility. The other candidates each introduce a vulnerability to nonextensibility or cause divergence.
@@
With dynamic scope, references resolve at the point of use rather than the definition site. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Dynamic scope breaks the $\lambda$-calculus embedding (Section 4). The Substitution Lemma (Lemma B.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect. With dynamic scope this separation fails. [^14] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$. The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$). With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$. With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$. This is the classical variable-capture failure of dynamic scoping. As with early binding, the mixin cannot serve as a complete replacement for functions.
@@
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (B.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix D). This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural. A closure captures a single environment, so a fixed-point combinator solves for a single self-reference. A stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix D). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
Conditional causes divergence
@@
The semantics of inheritance-calculus was not chosen from a menu. We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands. [^15] Three successive implementations confirmed this; see the discussion above. Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@ 7.1. Computational Models and $\lambda$-Calculus Semantics
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Semantically, the two share the same models: Meyer (Meyer, 1982) showed that combinatory algebras satisfying five additional conditions are exactly lambda algebras. No independent semantic domain for SKI has ever been exhibited from which $\lambda$-calculus semantics could be derived. The direction has always been the reverse: Scott’s $D_{\infty}$ was built for the $\lambda$-calculus, and SKI was interpreted inside it afterwards. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
@@
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991). It also requires a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem. Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975). Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions). In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies. None of these approaches is first-order; [^16] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics. none eliminates the function space from the semantic domain; and none unifies denotational and operational semantics into a single first-order definition. All three also inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
@@
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. These models eliminate Scott topology but not the function space itself: they remain higher-order, encoding $D\to D]$ inside $D$ via injections, intersection types, or multiset relations. Graph models (Engeler, 1981; Plotkin, 1993) encode functions via a web injection on powerset lattices; Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the set of intersection types derivable for it; Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the BCD filter model’s $\lambda$-theory coincides exactly with Böhm tree equality. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure. None is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds. None is first-order, since the function space is encoded, not eliminated. And none provides a fully abstract denotational semantics of the lazy $\lambda$-calculus without higher-order machinery.
Recursive equations and fixpoint semantics
Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational (SLD resolution), model-theoretic (least Herbrand model), and fixed-point (least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms). Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on. Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions; in inheritance-calculus, divergence already emerges as non-termination of recursive evaluation, so no separate coinductive formulation is needed. None of these prior frameworks has been applied to a first-order domain of paths to obtain a fully abstract semantics of the lazy $\lambda$-calculus without function spaces, substitution, or reduction rules. In particular, Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs, but has not been used to give recursive $\lambda$-programs a fixpoint semantics in which they converge instead of diverging.
7.2. Inheritance, Objects, and Records